### v2

@@ 1 Introduction
In this article, a new architecture called the Fast-Slow Neural Network (FSNN) is developed for learning data-driven models of singularly perturbed dynamical systems with dissipative fast timescale dynamics. We refer to our approach as an intelligent attractor since the method enforces the existence of a trainable, attracting invariant slow manifold [1] as a hard constraint. The constraint is imposed by way of a novel neural network architecture with structured weight matrices. By leveraging the neural ODE framework from [2], our technique produces deterministic evolution equations that approximately reproduce trajectory training data. When the training data comes from an unknown dynamical system, the method achieves data-driven full-order model discovery with stability guarantees due to provable existence of the slow manifold. In short, our architecture is, by design, a singularly perturbed dynamical system which contains an attractor.
@@
Nonlinear dimensionality reduction inspired by inertial manifold theory has a rich history (see [5] and the references therein) originating in the 1980s and 1990s [6, 4, 7, 8, 9, 10, 11, 12]. Data-driven approaches for discovering inertial manifolds in dynamical systems and using separations of scales to evolve a lower-dimensional reduced order model first materialized in the 90s [13, 14, 15, 16, 17, 18, 19] and recently reemerged in popularity due to boom in machine-learning-based modeling [20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32]. The work in this manuscript tackles a special class of systems with attracting invariant sets, namely fast-slow systems with dissipative fast timescale dynamics [1]. Unlike previous data-driven approaches, the FSNN is a full-order model which, using neural network architectures with special properties, strictly enforces the existence of an attracting slow manifold. This method may be viewed as a major step along a pathway in developing data-driven model discovery techniques with embedded invariant manifolds. In future work, we plan to address the general case where explicit timescale separation is not used as a crutch.
@@ 3 Fast-Slow Neural Networks (FSNNs)
The fast slow neural network (FSNN) is a transformation of an initial condition, $z_{0}$, and parameter, $\epsilon>0$, to a time series $\{z_{n}\}_{n=0}^{N}$ corresponding to times $t_{0},\dots,t_{N}$, where $z_{n}\in\mathbb{R}^{N}$ for $n=0,\dots,N$. The initial condition is transformed into fast-slow coordinates, $(y_{0},x_{0})$, via (7) where $h$ is represented using an invertible coupling flow network. A numerical integration is used to approximate the solution to the fast-slow system in (8) to obtain the time series $\{(y_{n},x_{n})\}_{n=0}^{N_{t}}$ for corresponding times $t_{0},\dots,t_{N}$, where $T$ is represented using a negative Schur form network, $B$ is represented using a low-rank bilinear map network, and $C$ and $g$ are represented using feed-forward neural networks. The trajectories are then transformed back into the original coordinate system using
(14) $z_{n}=h^{-1}\left(\left[\begin{array}[]{cc}y_{n}&x_{n}\end{array}\right]^{\rm T};\epsilon\right),\qquad n=1,\dots,N.$
The key components of Definition 3, including the invertible coupling flow network, Schur form network, and low-rank bilinear map network, will be defined in the following sections. First we will begin by stating the main result of the paper.
@@ 3.2 The Negative Schur Form Network
Figure 2 shows the regions of the $m{\rm-}p$ plane where both eigenvalues are less than zero (green region), greater than zero (red region), complex (blue region), and real (outside of the blue region). In order for ${\rm Re}(\lambda_{\pm})<0$, we must have $m<0$ and $p>0$. In order to enforce these constraints, we enforce the restrictions in (49). We remark that for the first constraint, $R\cos(\theta)<0$, $R$ and $\cos(\theta)$ must have opposite signs. If we focus on the region $\theta\in(-\pi/2,\pi/2)$ where $\cos(\theta)>0$, we would require $R<0$. If the region of $\theta$ were instead shifted by $\pi$, then we would require $R<0$. However, since $\mathcal{M}(R,r,\theta+\pi,\phi)=\mathcal{M}(-R,r,\theta,\phi)$, these choices are equivalent.
@@ 3.4 Numerical Integration
When the solution is initially on the slow manifold ($y=0$), the stages $Y^{i}$ for $i=1,\dots,s$ evaluate exactly to zero and therefore the numerical scheme can be simplified to an integration of the slow variables given by
@@ 5.1 Visualization of the Attracting Slow Manifold Property
A significant advantage of the FSNN is that, for any choice of model parameters, the network will have an attracting slow manifold that, due to Theorem 8, will attract solutions for sufficiently small values of $\epsilon$. The slow manifold is represented by the graph, $y=0$. In Figure 4, we demonstrate this property on multiple random initializations of FSNNs for $N_{y}=N_{x}=1$. In this figure, network a, b, and c are representing different FSNNs with different weights and therefore different slow manifolds. Multiple initial conditions are drawn randomly from a standard normal distribution (left plot) and are integrated using each FSNN to a later time $t=4$. The trajectory points at $t=4$ (blue dots) and the resulting slow manifold (orange curves) are plotted at this final time, demonstrating convergence of solutions to the slow manifolds inherent to each network.

### v3

@@ 1 Introduction
In this article, a new architecture called the Fast-Slow Neural Network (FSNN) is developed for learning data-driven models of singularly perturbed dynamical systems with dissipative fast timescale dynamics. The constraint is imposed by way of a novel neural network architecture with structured weight matrices. By leveraging the neural ODE framework from [1], our technique produces deterministic evolution equations that approximately reproduce trajectory training data. When the training data comes from an unknown dynamical system, the method achieves data-driven full-order model discovery with stability guarantees due to provable existence of the slow manifold. In short, our architecture is, by design, a singularly perturbed dynamical system which contains an attractor.
@@
Nonlinear dimensionality reduction inspired by inertial manifold theory has a rich history (see [4] and the references therein) originating in the 1980s and 1990s [5, 3, 6, 7, 8, 9, 10, 11]. Data-driven approaches for discovering inertial manifolds in dynamical systems and using separations of scales to evolve a lower-dimensional reduced order model first materialized in the 90s [12, 13, 14, 15, 16, 17, 18] and recently reemerged in popularity due to boom in machine-learning-based modeling [19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31]. The work in this manuscript tackles a special class of systems with attracting invariant sets, namely fast-slow systems with dissipative fast timescale dynamics [32]. Unlike previous data-driven approaches, the FSNN, a special form of neural ODE [1], is a full-order model which, using neural network architectures with special properties, strictly enforces the existence of an attracting slow manifold. This method may be viewed as a major step along a pathway in developing data-driven model discovery techniques with embedded invariant manifolds. In future work, we plan to address the general case where explicit timescale separation is not used as a crutch.
@@ 3 Fast-Slow Neural Networks (FSNNs)
A fast slow neural network (FSNN) is a special neural ODE that represents a pullback vector field on a phase space $Z\ni z$ of the form $V(z)=(Dh(z))^{-1}v(h(z))$, where $h$ is a parametric diffeomorphism and $v$ is a parametric vector field in the modified Fenichel normal form (8). The quantities $h,T,C,B,g$ are parameterized as follows. The diffeomorphism $h$ is represented as an invertible coupling flow network; $T$ is represented using a negative Schur form network; $B$ is represented using a low-rank bilinear map network; and $C,g$ are represented using multi-layer perceptrons.
Note that the corresponding flow maps $F_{\Delta t}$ for $V$ and $f_{\Delta t}$ for $v$ are related according to $F_{\Delta t}=h^{-1}\circ f_{\Delta t}\circ h$. It follows that computing the flow map $F_{\Delta t}$ from a FSNN reduces to first computing the flow map $f_{\Delta t}$ for $v$ and then appropriately applying $h$ or $h^{-1}$. The key components of Definition 3, including the invertible coupling flow network, Schur form network, and low-rank bilinear map network, will be defined in the following sections. First we will begin by stating the main result of the paper.
@@ 3.2 The Negative Schur Form Network
Figure 2 shows the regions of the $m{\rm-}p$ plane where both eigenvalues are less than zero (green region), greater than zero (red region), complex (blue region), and real (outside of the blue region). In order for ${\rm Re}(\lambda_{\pm})<0$, we must have $m<0$ and $p>0$. In order to enforce these constraints, we enforce the restrictions in (47). We remark that for the first constraint, $R\cos(\theta)<0$, $R$ and $\cos(\theta)$ must have opposite signs. If we focus on the region $\theta\in(-\pi/2,\pi/2)$ where $\cos(\theta)>0$, we would require $R<0$. If the region of $\theta$ were instead shifted by $\pi$, then we would require $R>0$. However, since $\mathcal{M}(R,r,\theta+\pi,\phi)=\mathcal{M}(-R,r,\theta,\phi)$, these choices are equivalent.
@@ 3.4 Numerical Integration
During the prediction of the learned system, one can use the above IMEX time integrator to integrate the full system, using small time steps to resolve the fast time scale. Alternatively, the proposed approach offers a unique capability to take larger time steps that are on the slow time scale when only the slow dynamics is of interest. When the solution is initially on the slow manifold ($y=0$), the stages $Y^{i}$ for $i=1,\dots,s$ evaluate exactly to zero and therefore the numerical scheme can be simplified to an integration of the slow variables given by
@@ 5.1 Visualization of the Attracting Slow Manifold Property
A significant advantage of the FSNN is that, for any choice of model parameters, the network will represent a fast-slow dynamical system with an attracting slow manifold that, due to Theorem 8, will attract solutions for sufficiently small values of $\epsilon$. The slow manifold is represented by the graph, $y=0$. Here we visualize the property of the FSNN without performing any training from dynamical system data. Instead, we demonstrate the property by randomly generating the trainable weights of FSNNs for a singularly perturbed dynamical system with $N_{y}=N_{x}=1$. The goal is to show that any initial condition for a randomly generated FSNN will naturally converge to its slow manifold. Note that for such a simple system, its slow manifold is known analytically.
In Figure 4, network a, b, and c are representing different FSNNs with different weights and therefore different slow manifolds. Multiple initial conditions are drawn randomly from a standard normal distribution (left plot) and are integrated using each FSNN to a later time $t=4$ corresponding to $\epsilon=0$. The trajectory points at $t=4$ (blue dots) and the resulting slow manifold (orange curves) are plotted at this final time, demonstrating convergence of solutions to the slow manifolds inherent to each network.